OpenStudy (anonymous):

if 10-x = y , then does 20-2x always equal 2y?

OpenStudy (anonymous):

Yes it is. You have multiplied the equation by 2. You can confirm this by drawing a graph of x against y.

OpenStudy (anonymous):

if you can divide 20-2x=2y to get 10-x=y, then it confirms that it will always be 2y. The same would apply to 30-3x=3y